A firm sells two goods (X and Y) that are related in consumption. The estimated demand and cost conditions are:

PX = 20 – 0.1QX – 0.005QY

PY = 70 0.3QY 0.1QX

MCX = 1 + 0.1QX

MCY = 2 + 0.25QY

What are the profit-maximizing levels of output for the two goods?

QX = 20, QY = 10

QX = 41, QY = 24

QX = 56, QY = 24

QX = 51, QY = 74

none of the above
Answer #1

Rx = Px.Qx = 20Qx - 0.1Qx2 - 0.005QyQx

Ry =  Py.Qy = 70Qy - 0.3Qy2 - 0.1QyQx

MRx = dRx/dQx = 20 - 0.2Qx - 0.005Qy

MRy = dRy/dQy = 70 - 0.6Qy - 0.1Qx

Profit is maximized where MRx = MCx and MRy = MCy

19 = 0.3Qx + 0.005Qy

68 = 0.85Qy + 0.1Qx

Qx = 62, Qy = 73 (none of the above)